OpenStudy (anonymous):

Carla is going to draw a marble from the bag shown below, replace it, and then draw another marble. What is the probability that Carla will get a green or yellow marble on the first draw and a blue on the second draw? (3 green, 4 yellow, 5 blue, 8 pink)

OpenStudy (anonymous):

\[\frac{7}{20}\times \frac{5}{20}\]

OpenStudy (anonymous):

green 3/20 yellow 1/5 blue 5/19

OpenStudy (anonymous):

probability of getting green OR yellow is \[\frac{7}{20}\] and the probability of getting blue is \[\frac{5}{20}=\frac{1}{4}\] and you want BOTH so multiply

OpenStudy (anonymous):

i get \[\frac{7}{80}\]

OpenStudy (anonymous):

you are finding the probability of getting blue on the second draw where one has already been taken out making the total 19.

OpenStudy (anonymous):

the probability of getting a blue on the first draw is in fact 1/4 but on the second draw it is 5/19